# Improving Referring Image Segmentation using Vision-Aware Text Features

### v1

@@ Abstract
Referring image segmentation is a challenging task that involves generating pixel-wise segmentation masks based on natural language descriptions. Existing methods have relied mostly on visual features to generate the segmentation masks while treating text features as supporting components. This over-reliance on visual features can lead to suboptimal results, especially in complex scenarios where text prompts are ambiguous or context-dependent. To overcome these challenges, we present a novel framework VATEX to improve referring image segmentation by enhancing object and context understanding with Vision-Aware Text Feature. Our method involves using CLIP to derive a CLIP Prior that integrates an object-centric visual heatmap with text description, which can be used as the initial query in DETR-based architecture for the segmentation task. Furthermore, by observing that there are multiple ways to describe an instance in an image, we enforce feature similarity between text variations referring to the same visual input by two components: a novel Contextual Multimodal Decoder that turns text embeddings into vision-aware text features, and a Meaning Consistency Constraint to ensure further the coherent and consistent interpretation of language expressions with the context understanding obtained from the image. Our method achieves a significant performance improvement on three benchmark datasets RefCOCO, RefCOCO+ and G-Ref. Code is available at: https://nero1342.github.io/VATEX_RIS
Keywords:
referring segmentation vision-aware text features CLIP localization multimodal understanding meaning consistency constraint
@@ 1 Introduction
Figure 1: VATEX outperforms previous state-of-the-art methods LAVT by utilizing vision-aware text features. Best view in zoom.
@@
Referring image segmentation (RIS) is an emerging new task in computer vision that predicts pixel-wise segmentation of visual objects in images from natural language cues. Compared to traditional segmentation  [3, 49, 42, 44, 51, 14, 47], RIS allows user-friendly interactions via text prompts to select and control the segmentation results, which is useful in various applications, especially interactive applications for image editing and robotics.
@@
A particular technique to solve RIS is to obtain a robust alignment between language and vision. Performing such an alignment presents significant challenges due to the nature of languages, which are highly ambiguous and context-dependent. Early alignment approaches [25, 53, 35, 11] in RIS either used bottom-up methods, merging vision and language features in early fusion and using an FCN as a decoder to produce object masks, or top-down methods, which first identify objects in the image and use the expression as the grounding criterion to select the best-matched result.
@@
Recent approaches [52, 9, 30] are based on transformers that learn the interaction between vision-text modalities followed by a standard encoder-decoder process to produce pixel-level segmentation results. However, this interaction still remains sufficiently underexplored in referring image segmentation. Specifically, the cross-modal interaction mainly occurs only in one way: keeping text features unchanged and utilizing them to enhance visual features. As a result, previous methods fail to identify the target object and generate precise segmentation for complex and challenging text expressions. For example, in Figure 1(a), although LAVT can easily segment the well-defined "orange-striped fish" and specific location information "on the left" in the first expression, LAVT struggles with the object "clownfish" and picks the turtle that is "facing towards the camera" in the second expression. We hypothesize that gradually putting the visual information into text features and performing the full bidirectional vision-text interactions can capture fine-grained semantic information of these modalities. This encourages us to focus more on vision-aware text features for object and context understanding.
First, for object understanding, we proposed integrating visual cues into text features in the query initialization process, facilitated by the visual-text embedding provided by CLIP [38]. We propose using the alignment between CLIP-Image and CLIP-Text embeddings to generate an object-centric visual heatmap that can be subsequently used as an initial mask for our model. By integrating this heatmap into the original text feature, we combine both linguistic and spatial information cues during the query initialization phase. This query initialization method transfers the knowledge from the pre-trained CLIP model and allows the model to localize the referred object correctly, even in the challenging case where the expression contains "unseen" category (e.g. clownfish).
@@
Second, for context understanding, we introduce a Contextual Multimodal Decoder (CMD) to further exploit the superior interaction between visual and text modalities, especially the vision-to-language interaction. CMD aims to enhance text features by using contextual information obtained from the visual features and to bring the semantic-aware textual information back to visual features in a hierarchical architecture. While we can use the ground truth mask annotations as a direct learning signal to supervise the language-to-vision features, the opposite interaction is implicitly learned without any learning signal. By observing that there are multiple ways to describe an instance based on the context provided by the image, we propose the Meaning Consistency Constraint to enforce consistency of vision-aware text features across different expressions referring to the same instance in an image. The vision-to-language interaction can explicitly learn through this extra in-context learning signal, resulting in a profound, coherent, and contextual understanding in the feature space.
We present our novel framework Vision-Aware TEXt for Referring Image Segmentation (VATEX), which achieves superior performance on all splits of the RefCOCO, RefCOCO+, and G-Ref datasets, surpassing the current state of the art for each dataset by $3-4\%$. We also achieve state-of-the-art performance on both the validation split of Ref-YouTube-VOS and Ref-DAVIS 2017 with $\mathbf{65.4}$ $\mathcal{J}\&\mathcal{F}$. Figure 1(b) demonstrates the differences in the complexity of the query sentences between three RIS benchmark datasets and the quantitative result of our proposed method compared to SOTA methods for RIS, which is LAVT [52] and VG-LAW [40]. In summary, our key contributions are as follows.
  • We propose a novel DETR-based framework for referring image segmentation, specially designed to utilize vision-aware text features.
  • We present a CLIP-based prior, a Contextual Multimodal Decoder, and a Meaning Consistency Constraint to address the complexities of cross-modal understanding and language ambiguities.
  • Through extensive experiments on three referring image segmentation benchmarks (RefCOCO, RefCOCO+, G-Ref) and two referring video segmentation benchmarks (Ref-YouTube-VOS, Ref-DAVIS17), our approach sets the new state-of-the-art.
@@ 2 Related Work
Referring image segmentation [15] aims to generate pixel-wise segmentation masks for referred objects in images given a natural language expression. Early works [29, 16, 12] proposed to extract visual and linguistic features independently from convolutional and recurrent neural networks, respectively, and then concatenating these features to create multimodal features for decoding final segmentation results. In recent works [24, 55, 52, 8, 10, 1], transformer-based multimodal encoders have been designed to fuse visual and text features, capturing the interaction between vision and language information in the early stage. VG-LAW [40] utilizes language-adaptive weights to dynamically adjust the visual backbone, enabling expression-specific feature extraction for better mask prediction while JMCELN [17] learns contextual embeddings and progressively aligns predictions for more accurate image segmentation. PolyFormer [30], on the other hand, treats this task as a sequential polygon generation process.
@@ 3 Proposed Method
Figure 2: The overall architecture of VATEX, which processes input images and language expressions through two concurrent pathways. Initially, the CLIP Prior module generates object queries, while simultaneously, traditional Visual and Text Encoders create multiscale visual feature maps and word-level text features. These visual and text features are passed into the Contextual Multimodal Decoder to enable multimodal interactions, yielding vision-aware text features and text-enhanced visual features. We then harness vision-aware text features to ensure semantic consistency across varied textual descriptions that reference the same object by employing sentence-level contrastive learning, as described in the Meaning Consistency Constraints section. On the other hand, the text-enhanced visual features and the object queries generated by the CLIP Prior are refined through a Masked-attention Transformer Decoder to produce the final output segmentation masks.
Our framework is constructed by three main components, as demonstrated in Figure 2. First, we propose a CLIP Prior module to generate an object-centric visual heatmap that localizes the object of interest from the text expression, which can be subsequently used to initialize the object queries for the DETR-based method (Section 3.1). Next, we propose to utilize cross-attention modules to interact between visual-text modalities in a hierarchical architecture via our Contextual Multimodal Decoder (Section 3.2). During training, we leverage Meaning Consistency Constraint to harness vision-aware text features generated by CMD to obtain a meaningful and consistent feature space (Section 3.3). We further adopt a masked-attention transformer decoder [6] to enhance the object queries through multiscale text-guided visual features. Finally, the enhanced object queries and the visual features from CMD are utilized to output segmentation masks (Section 3.4).
@@ 3.1 Object Localization with CLIP Prior
To accurately locate the object of interest in the image from a text expression, it is essential to identify the visual region that is the most closely related to the expression. In previous work, query-based methods typically convert only textual information from the natural language into object queries [46] or learn the target object representation implicitly through multi-modal transformers [8]. Our goal is to leverage the exceptional alignment between visual and text features in the CLIP [38] model to generate the heatmap explicitly and embed that positional prior in the query.
A naive approach of entering full text expressions and images into the CLIP Model results in poor localization accuracy due to the complex input sentence. CLIP is trained to map textual descriptions to visual features, enabling it to identify various objects and visual concepts. Due to the diversity of these visual concepts, only a few specific features are activated for a particular class or category while the majority of features remain inactive and become redundant features. As the complexity of text prompts increases, there is a corresponding increase of redundant features within the feature space. When calculating the similarity between an image and a text query, these redundant features also contribute to the final output, causing noisy activations in the heatmap and degrading the model’s performance, as illustrated in  Table 5. An example is provided in Figure 3 where the heatmap covers irrelevant regions in the image with spot activation and does not target to segment the main object of the sentence.
Additionally, the human approach to RIS does not involve parsing complex sentences entirely. Instead, we naturally break down a referring expression into its core components: the object of interest and its description with context information. Initially, the primary focus is on identifying what is the object mentioned in the expression (e.g. "the bull"). Following this, the search within the image is narrowed to objects that match the main object’s category. The final step involves using the specific characteristics or contextual information described in the expression to pinpoint the target object.
@@
Inspired by this, we first extract the main noun phrase from the expression (e.g. the bull) in order to obtain a simplified text expression. Referring expressions in the RIS task are object-centric, which means that the main noun phrase can be considered as the main object of the sentence. We then convert the complex referring expression to a simple template-based sentence before passing it to the CLIP Encoder. In our implementation, we use "A Photo of [Object]" as our template as it is the most common prompt to describe an object [54] in CLIP, where the resulting text feature is represented by $\mathcal{F}_{\text{text}}$. We found that this improves the accuracy of the heatmap in localizing the object of interest. In a separate flow, our input image goes through the CLIP Visual Encoder, resulting in features for multiple image tokens $\mathcal{F}_{\text{CLIP-Image}}\in\mathbb{R}^{\left(\frac{H}{16}\times\frac{W}{16}+1\right)\times C}$.
While this template approach can efficiently localize regions of interest, it may lead to information loss due to oversimplification (e.g. focusing only on the bull in this scenario). However, its primary function is to narrow down the search space by localizing a region containing the object of interest, not necessarily finding out the exact object. To find the exact object of interest, the full-text expression needs to pass through the CMD and MCC for more comprehensive characteristics and contextual understanding.
@@
As normal practice, positional prior from CLIP is embeded to the text features by changing the dimension of the similarity map from $\frac{H}{16}\times\frac{W}{16}+1$ to $C$, then repeat it $N$ times together with the text features, then add these two to create initial object queries feature with $N$ queries, each with $C$-dimension, for embedding the positional prior and text information into the query feature.
@@ 3.2 Contextual Multimodal Decoder
For a robust use of visual and text features in subsequent steps, we propose Contextual Multimodal Decoder (CMD) to produce multi-scale text-guided visual feature maps while enhancing contextual information from the image into word-level text features in a hierarchical design, see the architecture figure in the supplementary material. Specifically, our CMD is based on a feature pyramid network architecture [27], which has four levels. Each level transfers the semantic information from visual features to text features and then uses these vision-aware text features to update the visual features afterward via cross-modal attention.
@@
Figure 4: Illustration of Meaning Consistency Constraint. Vision-aware text embeddings of different expressions are passed through a contrastive learning module in sentence-level feature space. Embeddings referring to the same object are pulled closer while pushing others far away. The cosine similarities of text features before and after CMD are shown on the right. Best view in color.
@@ 3.3 Meaning Consistency Constraint
It is common for each object in an image to be described by different text expressions. Although the linguistic meaning of these descriptions may be different, they should convey the same semantic meaning when referencing that image (see two red expressions in  Figure 4). According to this perspective, it’s crucial for CMD to gradually comprehend contextual cues from visual features into textual representations and ensure consistent identification of target objects, where expressions referring to the same object yield identical representations. However, previous studies have often overlooked the relationship between expressions that pertain to the same instance.

### v2

# Vision-Aware Text Features in Referring Image Segmentation: From Object Understanding to Context Understanding

@@ Abstract
Referring image segmentation is a challenging task that involves generating pixel-wise segmentation masks based on natural language descriptions. The complexity of this task increases with the intricacy of the sentences provided. Existing methods have relied mostly on visual features to generate the segmentation masks while treating text features as supporting components. However, this under-utilization of text understanding limits the model’s capability to fully comprehend the given expressions. In this work, we propose a novel framework that specifically emphasizes object and context comprehension inspired by human cognitive processes through Vision-Aware Text Features. Firstly, we introduce a CLIP Prior module to localize the main object of interest and embed the object heatmap into the query initialization process. Secondly, we propose a combination of two components: Contextual Multimodal Decoder and Meaning Consistency Constraint, to further enhance the coherent and consistent interpretation of language cues with the contextual understanding obtained from the image. Our method achieves significant performance improvements on three benchmark datasets RefCOCO, RefCOCO+ and G-Ref. Project page: https://vatex.hkustvgd.com/.
@@ 1 Introduction
Referring image segmentation (RIS) is an emerging new task in computer vision that predicts pixel-wise segmentation of visual objects in images from natural language cues. Compared to traditional segmentation  [5, 57, 50, 52, 61, 16, 55], RIS allows users to select and control the segmentation results via text prompts, which is useful in various applications such as image editing, where users can modify specific parts of an image using simple text commands, and in robotics, where robots need to understand and act on verbal instructions in dynamic environments.
@@
A particular technique to solve RIS is to obtain a robust alignment between language and vision. Performing such an alignment presents significant challenges due to the nature of languages, which are highly ambiguous without given context. Early alignment approaches [31, 63, 41, 13] in RIS either used bottom-up methods, merging vision and language features in early fusion and using an FCN [46] as a decoder to produce object masks, or top-down methods, which first identify objects in image and use the expression as the grounding criterion to select best-matched result.
Figure 1: Qualitative comparison between LAVT and Ours. The yellow box indicates the wrong segmentation results. Object understanding and Context understanding are required to tackle the challenge of complex and ambiguous language expression.
@@
Recent approaches [62, 11, 36] are based on transformers that learn the interaction between vision-text modalities followed by a standard encoder-decoder process to produce pixel-level segmentation results. However, existing methods have relied mostly on visual features to generate the segmentation masks while treating text features as supporting components in the fusion module. This insufficient utilization of text understanding hampers these methods’ ability to accurately segment target objects for challenging expressions involving rare object vocabulary or contextual relationships between objects. For instance, as illustrated in  Figure 1, while LAVT [62] can effectively segment the well-defined ”orange-striped fish” and identify specific location information ”on the left” in the first expression, it struggles with the expression ”clownfish” and incorrectly identifies the turtle that is ”facing towards the camera” in the second expression. This inconsistency highlights the limitations of current approaches in understanding complex expressions, especially in handling unseen vocabulary and varying expressions referring to the same object.
Moreover, the human approach to RIS [2, 23, 1] does not involve parsing or understanding complex sentences entirely. Instead, we naturally break down a referring expression into its core components: the object of interest and its description with context information. Initially, the primary focus is on identifying what is the object mentioned in the expression (e.g., the main object of interest). Following this, the search space within the image is narrowed to objects that match the main object’s category. The final step involves using the specific characteristics or contextual information described in the expression to pinpoint the target object. Inspired by this, we propose decomposing this task into two processes: object understanding and context understanding. This decomposition allows for a more comprehensive understanding of text features, ultimately enhancing the accuracy and consistency of referring expression segmentation.
Firstly, in terms of object understanding, current methods do not utilize the object representation in the query initialization process. ReferFormer [54] generated object queries conditioned on language expressions, while VLT [10] implicitly employed multiple query vectors with different attention weights to generate various interpretations of the language description. However, these variations may lead to confusion and conflict with each other and may not focus on the target object. On the other hand, we propose CLIP Prior to explicitly integrate visual information of the primary object of interest into text cues during the query initialization process. This module transfers the knowledge from pre-trained model CLIP [44] and generates an object-centric visual heatmap to create adaptive, vision-aware queries, enhancing generalization and robustness of object comprehension, even in the challenging case where the expression contains ”unseen” category (e.g. clownfish).
@@
Secondly, for context understanding, we introduce a Contextual Multimodal Decoder (CMD) to further exploit the superior interaction between visual and text modalities, especially the vision-to-language interaction. CMD aims to enhance text features by using contextual information obtained from the visual features and to bring the semantic-aware textual information back to visual features in a hierarchical architecture. While we can use the ground truth mask annotations as a direct learning signal to supervise the language-to-vision features, the opposite interaction is implicitly learned without any learning signal. By observing that there are multiple ways to describe an instance based on the context provided by the image, we propose the Meaning Consistency Constraint (MCC) as a contrastive learning signal to enforce the consistency of vision-aware text features produced from CMD among different expressions referring to the same instance in an image. The vision-to-language interaction can explicitly learn through this extra in-context learning signal, resulting in a profound, coherent, and contextual understanding in the feature space.
Our method is evaluated on three widely-used image datasets, RefCOCO, RefCOCO+, and G-Ref, and further extends the results to video datasets, Ref-Youtube-VOS and Ref-DAVIS17. These datasets consist of diverse and challenging text expressions, and our proposed model achieves state-of-the-art performance on all five. Through various ablation studies, we have demonstrated the effectiveness of our model and shown that it can achieve robust referring segmentation even in challenging scenarios.
Our main contributions can be summarized as follows:
  • We address the current limitations of existing methods in dealing with complex text expressions and present a novel framework to utilize Vision-Aware TEXt Features (VATEX) for a better understanding of text expressions in RIS by decomposing it into Object Understanding and Context Understanding components.
  • We introduce a novel CLIP Prior to embed an object-centric visual heatmap in the query initialization process, enhancing object understanding by transferring knowledge from the pre-trained CLIP model.
  • We propose Contextual Multimodal Decoder (CMD) followed by a Meaning Consistency Constraint (MCC) as a learning signal for vision-to-language branch to improve context understanding. CMD module enhances the interaction between visual and text modalities, while MCC ensures consistent interpretation of different expressions conditioned in an image.
  • Our method achieves superior performance on all splits of the RefCOCO, RefCOCO+, and G-Ref for image datasets and Ref-YouTube-VOS and Ref-DAVIS 2017 for video datasets, surpassing the current state of the art for each dataset, especially in datasets with the more complex expressions.
Figure 2: The overall architecture of VATEX, which processes input images and language expressions through two concurrent pathways. Initially, the CLIP Prior module generates object queries, while simultaneously, traditional Visual and Text Encoders create multiscale visual feature maps and word-level text features. These visual and text features are passed into the Contextual Multimodal Decoder to enable multimodal interactions, yielding vision-aware text features and text-enhanced visual features. We then harness vision-aware text features to ensure semantic consistency across varied textual descriptions that reference the same object by employing sentence-level contrastive learning, as described in the Meaning Consistency Constraint section. On the other hand, the text-enhanced visual features and the object queries generated by the CLIP Prior are refined through a Masked-attention Transformer Decoder to produce the final output segmentation masks.
@@ 2 Related Work
Referring image segmentation [18] aims to generate pixel-wise segmentation masks for referred objects in images given a natural language expression. Early works [35, 19, 14] proposed to extract visual and linguistic features independently from convolutional and recurrent neural networks, respectively, and then concatenating these features to create multimodal features for decoding final segmentation results. In recent works [30, 66, 62, 10, 12, 3, 68, 51, 58, 64], transformer-based multimodal encoders have been designed to fuse visual and text features, capturing the interaction between vision and language information in the early stage. VG-LAW [47] utilizes language-adaptive weights to dynamically adjust the visual backbone, enabling expression-specific feature extraction for better mask prediction, while LISA [29] learns to generate segmentation masks based on Large Language Models. PolyFormer [36], on the other hand, treats this task as a sequential polygon generation. Another line of works have explored enhancing text understanding in RIS using graph-based methods: CMPC-RefSeg [20] classifies words into entity, attribute, relation, and other categories, building a graph with entities and attributes as nodes and relations as edges, while LSCM-RefSeg [22] constructs fully connected graph, then based on dependency parsing trees to prune unnecessary edges. In contrast, our model does not depend on graph convolutional networks, instead focusing on utilizing the vision-aware text features in text understanding.
@@ 3 Proposed Method
Inspired by the human approach to RIS, which involves breaking down a referring expression into its core components: object of interest and contextual description, we propose simplifying the text expression by decomposing it into object and context parts. This decomposition aims to enhance the text understanding, thus improving the accuracy and consistency of referring expression segmentation.
Our framework is constructed by three main components, as demonstrated in Figure 2. First, for object understanding, we propose a CLIP Prior module to generate an object-centric visual heatmap that localizes the object of interest from the text expression, which can be subsequently used to initialize the object queries for the DETR-based method (Section 3.1). Next, we utilize cross-attention modules to interact between visual-text modalities in a hierarchical architecture via our Contextual Multimodal Decoder (Section 3.2) and leverage Meaning Consistency Constraint to harness vision-aware text features generated by CMD (Section 3.3). We further adopt a masked-attention transformer decoder [8] to enhance the object queries through multiscale text-guided visual features. Finally, the enhanced object queries and the visual features from CMD are utilized to output segmentation masks (Section 3.4).
@@ 3.1 Object Localization with CLIP Prior
For the object part, we first extract the main noun phrase from the expression using spaCy [17] (e.g., the bull) in order to focus only on this main noun phrase. Referring expressions in the RIS task are object-centric, which means that the main noun phrase is the main object of the sentence. We then convert the complex referring expression to a simple template-based sentence before passing it to the CLIP Encoder. In our implementation, we use ”A Photo of [Object]” as our template as it is the most common prompt to describe an object [65] in CLIP, where the resulting text feature is represented by $\mathcal{F}_{\text{text}}$. We found that this improves the accuracy of the heatmap in localizing the object of interest. In a separate flow, our input image goes through the CLIP Visual Encoder, resulting in features for multiple image tokens $\mathcal{F}_{\text{CLIP-Image}}\in\mathbb{R}^{\left(\frac{H}{16}\times\frac{W}{16}+1\right)\times C}$.
@@
As normal practice, positional prior from CLIP is embeded to the text features by changing the dimension of the similarity map from $\frac{H}{16}\times\frac{W}{16}+1$ to $C$, then repeat it $N$ times together with the text features, then add these two to create initial object queries feature with $N$ queries, each with $C$-dimension, for embedding the positional prior and text information into the query feature. Unlike previous methods that typically learn the target object representation implicitly through multi-modal transformers [10] or convert only textual information from natural language into object queries [54], our approach explicitly generates the heatmap and embeds it in the query initialization process. This ensures that the queries contain rich and useful information about both visual and textual aspects, as well as the alignment between these modalities. Such a comprehensive query initialization is essential for effective object understanding, particularly in challenging scenarios involving complex or unseen vocabulary.
While this template approach can efficiently localize regions of interest, it may lead to information loss due to oversimplification (e.g., focusing only on the bull in this scenario). However, its primary function is to narrow down the search space by localizing a region containing the object of interest, not necessarily finding out the exact object. To find the exact object of interest, the full-text expression needs to pass through the CMD and MCC for more comprehensive characteristics and contextual understanding.
@@ 3.2 Contextual Multimodal Decoder
Contextual Multimodal Decoder (CMD) is proposed to produce multi-scale text-guided visual feature maps while enhancing contextual information from the image into word-level text features in a hierarchical design, see the architecture figure in the supplementary material. Specifically, our CMD is based on a feature pyramid network architecture [33], which has four levels. Each level transfers the semantic information from visual features to text features and then uses these vision-aware text features to update the visual features afterward via cross-modal attention.
@@
Previous methods have developed various bidirectional multimodal fusion mechanisms, including word-pixel alignment in encoder stage [58, 64] and region-language interactions [34]. Compared to these approaches, our novelty is the combination of CMD and MCC, where MCC serves as an in-context learning signal to enrich the vision-aware text features and further enhance the vision-language interactions within the hierarchical architecture of CMD.
Method Backbone RefCOCO RefCOCO+ G-Ref
Visual Textual val testA testB val testA testB val test
VLT [11] Swin-B BERT 72.96 75.96 69.60 63.53 68.43 56.92 63.49 66.22
LAVT [62] Swin-B BERT 74.46 76.89 70.94 65.81 70.97 59.23 63.34 63.62
LISA-7B [29] ViT-H SAM Vicuna-7B 74.10 76.50 71.10 65.10 67.40 56.50 66.40 68.50
VG-LAW [47] ViT-B BERT 75.05 77.36 71.69 66.61 70.30 58.14 65.36 65.13
Methods Pr@0.5 Pr@0.7 Pr@0.9 mIoU
LAVT [62] 84.46 75.28 34.30 74.46
ReLA [34] 85.92 77.71 34.99 75.61
CG-Former [48] 87.23 78.69 38.77 76.93
VATEX (Ours) 88.12 82.54 45.11 78.17
Table 3: Quantitative results on video datasets.
Methods Ref-YT-VOS Ref-DAVIS17
[…]
ReferFormer [54] 62.9 61.3 67.5 61.1 58.1 64.1
VLT  [11] 63.8 61.9 65.6 61.6 58.9 64.3
VATEX (Ours) 65.4 63.3 67.5 65.4 62.3 68.5
@@ 3.3 Meaning Consistency Constraint
Figure 4: Illustration of Meaning Consistency Constraint. Vision-aware text embeddings of different expressions are passed through a contrastive learning module in sentence-level feature space. Embeddings referring to the same object are pulled closer while pushing others far away. Best view in color.
Each object in an image can be described by various text expressions. Although the linguistic meaning of these descriptions may be different, they should convey the same semantic meaning when referencing that image (see two red expressions in  Figure 4). According to this perspective, it’s crucial for CMD to gradually comprehend contextual cues from visual features into textual representations and ensure consistent identification of target objects, where expressions referring to the same object yield identical representations. However, previous studies have often overlooked the relationship between expressions that pertain to the same instance.